# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

### v3

@@ 4.1.3. Codebook Construction and Optimization
The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (Van Den Oord et al., 2017). During training, the continuous latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$ are quantized into discrete space by mapping each vector to its nearest codebook entry, $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$, based on minimizing the Euclidean distance. This process translates the encoded features into discrete tokens, enabling the model to leverage a finite set of vectors to capture complex patterns.
@@
To encourage equal usage of the codebook vectors (Baevski et al., 2020), we incorporate a diversity loss to ensure balanced representations across the codebook. It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.1cm}$
@@
However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
(5) $\mathcal{L}_{\text{ortho}}=\frac{1}{K^{2}}\left|\left|\ell_{2}(\mathbf{e})^{\top}\ell_{2}(\mathbf{e})-I_{K}\right|\right|_{F}^{2}\text{,}\vspace{-0.1cm}$
@@ 4.2. TS and CS Decoders
(6) $\mathcal{L}_{1}=$ $\lambda_{\text{ortho}}\mathcal{L}_{\text{ortho}}+\lambda_{\text{div}}\mathcal{L}_{\text{div}}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{ts}\right|\right|^{2}_{2}+\left|\left|\mathbf{x}-\mathbf{x}^{\prime}_{cs}\right|\right|^{2}_{2}$
$+$ $\left|\left|sg[\mathbf{z}_{e}^{ts}(\mathbf{x})]-\mathbf{z}_{q}^{ts}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{ts}(\mathbf{x})]-\mathbf{z}_{e}^{ts}(\mathbf{x})\right|\right|^{2}_{2}$
$+$ $\left|\left|sg[\mathbf{z}_{e}^{cs}(\mathbf{x})]-\mathbf{z}_{q}^{cs}(\mathbf{x})\right|\right|^{2}_{2}+\left|\left|sg[\mathbf{z}_{q}^{cs}(\mathbf{x})]-\mathbf{z}_{e}^{cs}(\mathbf{x})\right|\right|^{2}_{2}\text{,}$
@@ 4.3.1. Feature Fusion and Alignment
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021). The architecture incorporates cross-attention layers, balancing both accuracy and computational efficiency, ensuring that factors derived from the combined features are more representative of market dynamics.
Contrastive Learning. To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (Li et al., 2023; Radford et al., 2021). Within the factor module, a contrastive loss encourages similar features to cluster closely while pushing apart dissimilar ones in the embedding space.
@@ 4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (Dosovitskiy et al., 2020), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (Duan et al., 2022). Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
(7) $[\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
@@
(8) $[\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
@@ 4.4. Downstream Tasks
Existing research on latent factor models (Duan et al., 2022; Wei et al., 2023) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in quantitative trading is typically evaluated through portfolio management tasks. In contrast, STORM captures both the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
@@ 4.4.1. Portfolio Management
To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (Duan et al., 2022), focusing on return prediction and the PM task. Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html to backtest the factor model. The TopK-Drop strategy constructs a daily portfolio consisting of top $k$ stocks based on the return predictions. Considering the transaction costs in the real stock market, the strategy sets the turnover constraint, which ensures that more than $k-d$ stocks overlap between the portfolios $\mathbf{w}_{t}$ and $\mathbf{w}_{t+1}$ on consecutive trading days.
@@ 4.4.2. Algorithmic trading
In STORM, $\mathbf{z}_{e}(\mathbf{x})$ is the stock latent factor embeddings encoded by the dual VQ-VAE architecture, including both time-series and cross-sectional factor embeddings in the data. The latent factor embeddings $\mathbf{Z}$ are integrated into the observation set $\mathcal{O}=\{\mathbf{Z},\mathcal{R}\}$, where $\mathcal{R}$ is the reward function used to guide the agent’s learning and decision-making in the environment. By combining latent factor embeddings with reward, the agent can more accurately understand the market environment and adjust its strategy based on observations. In the AT task, since it involves trading a single stock, the features of other stocks are excluded. We use the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) to optimize the policy.
@@ 5. Experiment
In this section, we evaluate the proposed STORM on real stock markets and conduct extensive experiments to address the following research questions. RQ1: How to evaluate the effectiveness of STORM’s learned factors? RQ2: How does STORM perform on downstream tasks? RQ3: How do the key components contribute to the performance of STORM?
@@ 5.1. Experiment Settings
Dataset. We conduct experiments on two U.S. stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (Yang et al., 2020). Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19. Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
Metrics. We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks. The financial metrics include 2 profit criteria, Annualized Percentage Yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar Ratio (CR) and Annualized Sharpe Ratio (ASR), and 2 risk criteria, Maximum Drawdown (MDD) and Annualized Volatility (AVO). The calculation formulas and meanings of these metrics are as follows:
@@
    (10) $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
@@
  • RankICIR is the information ratio of RankIC, which measures the stability of prediction, $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}$.
Baselines. As shown in Table 2, we compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods can be divided into three categories: ML & DL-based models, Factor models, and RL-based models. We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task. Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks. The following will provide a brief introduction to each method:
  • Market
  • ML&DL-based
    • LGBM (Yang et al., 2020) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
    • LSTM (Yu and Li, 2018) utilizes long short-term memory to improve the accuracy of price predictions.
    • Transformer (Yang et al., 2020) models leverage self-attention mechanisms to enhance the precision of price forecasts.
  • RL-based
    • SAC (Haarnoja et al., 2018) is an off-policy actor-critic algorithm that optimizes trading strategies using entropy regularization and soft value functions in continuous action spaces.
    • PPO (Schulman et al., 2017) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and sample efficiency.
    • DQN (Mnih et al., 2013) uses deep neural networks to approximate the action-value function and make trading decisions from market data.
  • Factor-model
    • CAFactor (Gu et al., 2021) is a state-of-the-art model that learns latent factors depending on asset characteristics.
    • FactorVAE (Duan et al., 2022) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
    • HireVAE (Wei et al., 2023) is a state-of-the-art VAE-based latent factor model that identifies factors between market situation and stock-wise.
Models SP500 DJ30
[…]
Future Return Prediction in Portfolio Management Task
LightGBM 0.027 0.274 0.031 0.272
± 0.006 ± 0.084 ± 0.005 ± 0.049
LSTM 0.034 0.333 0.031 0.329
± 0.006 ± 0.042 ± 0.004 ± 0.056
Transformer 0.035 0.340 0.033 0.343
± 0.007 ± 0.078 ± 0.005 ± 0.045
CAFactor 0.037 0.356 0.040 0.380
± 0.005 ± 0.084 ± 0.003 ± 0.043
FactorVAE 0.052 0.543 0.056 0.520
± 0.010 ± 0.122 ± 0.012 ± 0.081
HireVAE 0.057 0.558 0.058 0.563
± 0.006 ± 0.058 ± 0.006 ± 0.053
STORM 0.063 0.679 0.066 0.670
± 0.020 ± 0.169 ± 0.038 ± 0.303
STORM-w/o-TS 0.051 0.514 0.055 0.560
± 0.016 ± 0.141 ± 0.016 ± 0.105
STORM-w/o-CS 0.053 0.514 0.053 0.586
± 0.013 ± 0.126 ± 0.017 ± 0.143
Improvement(%)11footnotemark: 1 10.526 21.685 13.793 19.005
  • $1$
    Improvement of STORM over the best-performing baselines.
Table 1. Results of all models on two metrics, i.e., RankIC and RankICIR (mean ± range, computed across 10 runs).
Task Market ML & DL-based RL-based Factor model Ours
B&H LGBM (Yang et al., 2020) LSTM (Yu and Li, 2018) Transformer (Yang et al., 2020) SAC (Haarnoja et al., 2018) PPO (Schulman et al., 2017) DQN (Mnih et al., 2013) CAFactor (Gu et al., 2021) FactorVAE (Duan et al., 2022) HireVAE (Wei et al., 2023) STORM
Prediction
Portfolio Management
Algorithmic Trading
Table 2. Comparison of baselines across different tasks
Implementation Details. All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each. The implementations of the ML & DL methods are based on Qlib (Yang et al., 2020). The input data is organized into batches with dimensions corresponding to historical days, stocks, and features, ensuring compatibility across tasks. As for other baselines, we adopt the default settings in their public implementations. Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported. Detailed information regarding the network architecture of STORM, as well as the optimizers and hyperparameters used for the two downstream tasks, is provided in Appendix B.
@@
STORM 0.177 1.641 1.193 0.817 0.692 0.072
± 0.103 ± 0.457 ± 0.280 ± 0.071 ± 0.091 ± 0.011
STORM-w/o-TS 0.084 1.278 1.055 0.702 0.632 0.068
± 0.060 ± 0.204 ± 0.271 ± 0.062 ± 0.107 ± 0.003
STORM-w/o-CS 0.080 1.264 0.971 0.698 0.684 0.068
± 0.052 ± 0.176 ± 0.192 ± 0.065 ± 0.071 ± 0.005
Improvement(%) 124.05 30.65 195.30 77.61 - 60.47
@@
STORM 0.143 1.501 0.839 0.680 0.644 0.057
± 0.078 ± 0.321 ± 0.213 ± 0.080 ± 0.102 ± 0.004
STORM-w/o-TS 0.071 1.244 0.757 0.596 0.634 0.056
± 0.134 ± 0.509 ± 0.375 ± 0.144 ± 0.159 ± 0.004
STORM-w/o-CS 0.077 1.266 0.764 0.607 0.632 0.057
± 0.156 ± 0.612 ± 0.203 ± 0.18 ± 0.118 ± 0.004
Improvement(%) 88.16 20.47 138.35 41.67 - 63.40
@@
5.2. Factor Quality (for RQ1)
To demonstrate the effectiveness and quality of the factors, we designed a stock future return prediction task to showcase the factors’ predictive ability and robustness to capture dynamic market changes. In the prediction task, STORM shows substantial improvements over six baseline methods across the SP500 and DJ30 datasets. Specifically, in terms of RankIC and RankICIR, STORM achieves an average improvement of 16.105% on SP500 and 16.399% on DJ30, as shown in Table  1. These results highlight the learned latent factors are significantly effective in capturing stock price trends.
The diversity of factors is essential in financial investment, as it allows the model to capture the multi-dimensional characteristics of the market, avoiding redundancy or over-reliance on a single factor. In the codebook, each embedding represents a cluster center for a category of factor embeddings. To analyze the utilization rate of each factor category, we report the frequency of codebook embedding indices in the testing data. For clarity, we group every five indices into one interval, as illustrated in Figure 3(a). The results indicates a relatively even distribution of factor usage, highlighting the model’s effectiveness to maintain diversity during the factor selection process.
Furthermore, we experiment with different codebook sizes on the DJ30 dataset to find the optimal balance between factor diversity and representation quality. A larger codebook may lead to sparse factor representations, reducing the effectiveness of factor utilization, while a smaller codebook could result in insufficient factor differentiation and information redundancy. Following standard practices, we evaluate five different codebook sizes, including 128, 256, 512, 1024, and 2048. As shown in Figure 3(b), codebook size of 512 achieves optimal factor prediction performance, striking a balance between expressiveness and efficiency.
Figure 3. (a) Distribution of counts for codebook embedding indices. (b) the RankIC and RankICIR results across different codebook sizes.
5.3. Downstream Task Performance (for RQ2)

### v4

@@ 4.1.3. Codebook Construction and Optimization
The construction of the codebook involves a learnable embedding space designed to quantize the encoded features into discrete representations (van2017neural). During training, the continuous latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$ are quantized into discrete space by mapping each vector to its nearest codebook entry, $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$, based on minimizing the Euclidean distance. This process translates the encoded features into discrete tokens, enabling the model to leverage a finite set of vectors to capture complex patterns.
@@
To encourage equal usage of the codebook vectors (baevski2020wav2vec), we incorporate a diversity loss to ensure balanced representations across the codebook. It is achieved by maximizing the entropy of the averaged softmax distribution over vectors for each codebook $\bar{p}_{g}$:
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}$
@@
However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (shin2023exploration) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
[…]
@@ 4.2. TS and CS Decoders
[…]
@@ 4.3.1. Feature Fusion and Alignment
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two temporal and spatial patterns and produce stronger fusion features for factor generations, motivated by CrossViT (chen2021crossvit). The architecture incorporates cross-attention layers that explicitly learn correlations between two patterns, enabling the encoder to emphasize mutually informative regions while filtering out noise. This design balances accuracy and computational efficiency, ensuring the fused features yield richer and more discriminative signals for factor generation and that the resulting factors are more representative of market dynamics.
Contrastive Learning. To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (li2023blip; radford2021learning). A contrastive loss is introduced within the factor module to encourage features from TS and CS modules that represent similar market conditions to cluster closely in the embedding space, while pushing apart features corresponding to different conditions. This dual-stream alignment improves the coherence of the factor space and enhances the model’s ability to generalize across diverse market scenarios.
@@ 4.3.2. Prior-Posterior Learning
Through multi-scale encoder and contrastive learning layer, the CS and TS latent features are fused and aligned. In order to retain the codebook’s categorical properties for factors, we incorporate the codebook embeddings $\mathbf{z}_{q}^{ts}(\mathbf{x})$ and $\mathbf{z}_{q}^{cs}(\mathbf{x})$ as extra [CLS]-tokens (dosovitskiy2020image), which are appended to the latent features $\mathbf{z}_{e}^{ts}(\mathbf{x})$ and $\mathbf{z}_{e}^{cs}(\mathbf{x})$, respectively. Then, two latent features concat together to get the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$.
Given the inherent volatility and complexity of the stock market, bridging the gap between noisy market data and an effective factor model remains a significant challenge (duan2022factorvae). Therefore, we introduce the prior-posterior structure to predict future returns and optimize latent factor embeddings. In the training stage, the posterior layer estimates posterior distribution of factor expected returns $\mathbf{z}_{post}$ from the true future stock returns $\mathbf{y}$ and the latent factor embeddings $\mathbf{z}_{e}(\mathbf{x})$:
(7) $\mu_{post},\sigma_{post}]=\phi_{FE}(\mathbf{y},\mathbf{z}_{e}(\mathbf{x}))\text{,}$
@@
(8) $\mu_{prior},\sigma_{prior}]=\phi_{FP}(\mathbf{z}_{e}(\mathbf{x}))\text{.}$
@@ 4.4. Downstream Tasks
Existing research on latent factor models (duan2022factorvae; wei2023hirevae) primarily focuses on predicting future returns by analyzing cross-sectional characteristics of market stocks, which means the profitability of the factor model in financial investment is typically evaluated through PM tasks. In contrast, STORM captures the cross-sectional features among stocks and the time-series features of individual stocks, enabling it to excel in both PM and single-asset AT tasks.
@@ 4.4.1. Portfolio Management
To evaluate the performance of the STORM and compare it with baseline methods, we follow the paradigm proposed by FactorVAE (duan2022factorvae), focusing on return prediction and the PM task. Specifically, we utilize the factor decoder network to generate stock future returns $\mathbf{\hat{y}}$, and then apply the TopK-Drop strategy [^3] https://qlib.readthedocs.io/en/latest/component/strategy.html, which constructs a daily portfolio by selecting the top $k$ stocks based on predicted returns, to backtest the factor model. Considering transaction costs in real markets, there is a turnover constraint that limits the number of changes in the portfolio composition between consecutive trading days. Specifically, when moving from $\mathbf{w}_{t}$ to $\mathbf{w}_{t+1}$, at most $d$ stocks can be replaced, where a replacement is defined as a stock whose weight changes from 0 (not held) to a positive value (newly bought) or from a positive value to 0 (fully sold). This constraint ensures that at least kdk-d stocks remain in the portfolio, reducing excessive turnover.
@@ 4.4.2. Algorithmic trading
In STORM, $\mathbf{z}_{e}(\mathbf{x})$ is the stock latent factor embeddings encoded by the dual VQ-VAE architecture, including both time-series and cross-sectional factor embeddings in the data. The latent factor embeddings $\mathbf{Z}$ are integrated into the observation set $\mathcal{O}=\{\mathbf{Z},\mathcal{R}\}$, where $\mathcal{R}$ is the reward function used to guide the agent’s learning and decision-making in the environment. By combining latent factor embeddings with reward, the agent can more accurately understand the market environment and adjust its strategy based on observations. In the AT task, since it involves trading a single stock, the features of other stocks are excluded. We use the Proximal Policy Optimization (PPO) algorithm (schulman2017proximal) to optimize the policy. In each iteration, the agent interacts with the environment to collect trajectories, computes advantage estimates, and updates the policy parameters via PPO’s clipped surrogate objective. This optimization process ensures the stability and efficiency of policy updates, helping the agent gradually improve the decision-making ability and achieve better trading strategies.
Table 1. Comparison of baselines across different tasks
Task Market ML & DL-based RL-based Factor model Ours
B&H LGBM (yang2020qlib) LSTM (yu2018forecasting) Transformer (yang2020qlib) SAC (haarnoja2018soft) PPO (schulman2017proximal) DQN (mnih2013playing) CAFactor (gu2021autoencoder) FactorVAE (duan2022factorvae) HireVAE (wei2023hirevae) STORM
Prediction
Portfolio Management
Algorithmic Trading
@@ 5. Experiment
In this section, we evaluate the proposed STORM on real stock markets and conduct extensive experiments to address the following research questions.
RQ1:
How does STORM perform on downstream tasks?
RQ2:
How to evaluate the effectiveness of STORM’s learned factors?
RQ3:
How do the key components contribute to the performance of STORM?
@@ 5.1. Experiment Settings
5.1.1. Datasets
We conduct experiments on two U.S. stock markets, SP500 and DJ30, using stock daily data that includes technical features based on Alpha158 (yang2020qlib). Both datasets span 16 years, from 2008-04-01 to 2024-03-31, encompassing global conditions, e.g., the 2007-2008 financial crisis and COVID-19. Datasets are chronologically divided into non-overlapping training (from 2008-04-01 to 2021-03-31) and test (from 2021-04-01 to 2024-03-31) sets.
5.1.2. Metrics
We compare STORM and baselines in terms of 6 financial metrics across the PM and AT tasks. The financial metrics include 2 profit criteria, annualized percentage yield (APY) and Cumulative Wealth (CW), 2 risk-adjusted profit criteria, Calmar ratio (CR) and annualized Sharpe ratio (ASR), and 2 risk criteria, maximum drawdown (MDD) and annualized volatility (AVO). The calculation formulas and meanings of these metrics are as follows:
@@ 5.1.2. Metrics
    $\text{RankIC}_{s}=\frac{1}{N}\frac{(r_{\hat{y}_{s}}-\text{mean}(r_{\hat{y}_{s}}))^{T}(r_{y_{t}}-\text{mean}(r_{y_{s}}))}{\text{std}(r_{\hat{y}_{s}})\cdot\text{std}(r_{y_{s}})}$
@@
  • RankICIR is the information ratio of RankIC, which measures the stability of prediction,
    $\text{RankICIR}=\frac{\text{mean}(\text{RankIC}_{s})}{\text{std}(\text{RankIC}_{s})}\text{.}$
@@
STORM 0.18822footnotemark: 2 1.683 1.189 1.052 0.166 0.171
± 0.055 ± 0.226 ± 0.661 ± 0.329 ± 0.050 ± 0.020
STORM-w/o-TS 0.090 1.300 0.503 0.570 0.181 0.175
± 0.074 ± 0.277 ± 0.390 ± 0.339 ± 0.043 ± 0.025
STORM-w/o-CS 0.089 1.294 0.623 0.592 0.146 0.167
± 0.045 ± 0.163 ± 0.385 ± 0.252 ± 0.073 ± 0.020
Improvement(%)33footnotemark: 3 137.97 34.00 194.31 128.70 26.26 2.91
@@
STORM 0.148 1.517 1.396 1.052 0.108 0.140
± 0.046 ± 0.188 ± 0.679 ± 0.297 ± 0.026 ± 0.014
STORM-w/o-TS 0.079 1.259 0.621 0.603 0.127 0.142
± 0.057 ± 0.192 ± 0.425 ± 0.379 ± 0.038 ± 0.008
STORM-w/o-CS 0.073 1.236 0.533 0.566 0.138 0.140
± 0.051 ± 0.169 ± 0.400 ± 0.343 ± 0.034 ± 0.017
Improvement(%) 94.74 21.75 296.59 119.17 52.00 8.50
  • 11
    Underline indicates the best-performing baseline method result;
  • 22
    Bold indicates the best performance among our proposed models (including variants);
  • 33
    Improvement of STORM over the best-performing baselines.
@@
5.1.3. Baselines
As shown in Table 1, we compare STORM with nine methods across one prediction task and two downstream tasks. The baseline methods fall into three categories: ML & DL-based models, Factor models, and RL-based models. We evaluate STORM against ML&DL models and factor models in prediction and PM tasks, and against ML&DL models and RL methods in the AT task. Moreover, STORM is a versatile all-rounder that supports three tasks, unlike other methods specialized in specific tasks.
  • Market
  • ML&DL-based
    • LGBM (yang2020qlib) uses a series of tree models to predict price fluctuations and provide buy and sell signals.
    • LSTM (yu2018forecasting) utilizes long short-term memory to improve the accuracy of price predictions.
    • Transformer (yang2020qlib) models leverage self-attention mechanisms to enhance the precision of price forecasts.
  • RL-based
    • SAC (haarnoja2018soft) is an off-policy actor-critic algorithm using entropy regularization and soft value functions.
    • PPO (schulman2017proximal) updates trading policies iteratively to balance exploration and exploitation, ensuring stability and efficiency.
    • DQN (mnih2013playing) approximates the action-value function and make trading decisions from market data with a deep Q-network.
  • Factor-model
    • CAFactor (gu2021autoencoder) learns latent factors from asset features.
    • FactorVAE (duan2022factorvae) is a state-of-the-art VAE-based latent factor model that learns optimal latent stock factors.
    • HireVAE (wei2023hirevae) is a state-of-the-art VAE-based latent factor model capturing market and inter-stock factors.
5.1.4. Implementation Details
All experiments are implemented using PyTorch and conducted on two NVIDIA RTX H100 GPUs with 80GB of memory each. The input data is organized into batches with dimensions $(B,W,N,D)$, where $B=16$ is the batch size, $N$ is the number of stocks (28 for DJ30, 408 for SP500), $W=64$ denotes historical days, and $D=152$ is the total number of features, including the OHLC data and technical indicators.
The TS module uses patches of size $(4,1,D)$ (4-day sequences per stock), whereas the CS module uses $(1,N,D)$ (all stocks on a given day). Encoders (4 layers, 4 heads) and decoders (2 layers, 8 heads) are standard Transformer blocks with a shared embedding size of 256. Codebook sizes are searched in 128,256,512,1024,2048{128,256,512,1024,2048}, with 512 performing best.
[…]
For downstream tasks, we use Qlib’s [^4] https://github.com/microsoft/qlib TopK-Drop portfolio strategy (k=5k{=}5, d=3d{=}3) and PPO for trading. The pretrained TS/CS factors serve as states; all CS factors are retained while only TS factors relevant to traded stocks are used. The action space is buy, hold, sell. Rewards follow $r=\frac{v_{\text{post}}-v_{\text{pre}}}{v_{\text{pre}}}$ with $v=p\times m+c$, starting from $10^{6}$ initial cash and transaction cost $10^{-4}$.
The implementations of the ML & DL methods are based on Qlib (yang2020qlib). As for other baselines, we adopt the default settings in their public implementations. Each experiment is repeated multiple times with different random seeds, and the mean and variance of the metrics are reported.
5.2. Performance on Investment Tasks (for RQ1)